### v1

# Quantum-Enhanced LLM Efficient Fine Tuning

Abstract
Low-Rank Adaptation (LoRA) enables efficient fine-tuning of pre-trained language models via low-rank matrix approximation, which is effective in many scenarios. However, its low-rank representation capacity is constrained in complex tasks or high-rank dependency settings, potentially limiting model adaptability. Addressing the expressive bottleneck of classical low-rank approximation in fine-tuning large language models, this paper proposes a parameter-efficient fine-tuning method based on a Quantum Weighted Tensor Hybrid Network (QWTHN), which leverages Quantum Neural Network (QNN). The study investigates quantum-classical hybrid parameter-efficient fine-tuning in low-rank spaces. QWTHN decomposes pre-trained weights into quantum neural network and tensor network representations, utilizing quantum state superposition and other methods to break through classical rank limitations. Experiments show that the proposed quantum fine-tuning technique for large models approaches or even surpasses the parameter efficiency of LoRA. On the CPsyCounD and R1-Distill-SFT datasets, QWTHN, compared to classical LoRA, reduces training loss by up to 15% while using 76% fewer parameters, and achieves an 8.4% performance improvement on the CPsyCounD test set. This research not only realizes lightweight and efficient adaptation of quantum resources to billion-parameter models but also validates the practical path of quantum hardware driven by large model tasks, laying the first engineering-ready technical foundation for future quantum-enhanced AGI systems.
Index Terms:
Quantum Computing; LLM; Quantum LLM; Fine Tuning
I Introduction
The rapid advancement of large language models (LLMs) has driven innovations in Parameter-Efficient Fine-Tuning (PEFT) to reduce computational overhead while preserving performance. Classical methods such as Low-Rank Adaptation (LoRA) [1] and DoRA [2] operate under the assumption that weight updates during fine-tuning reside in low-rank subspaces, enabling efficient adaptation through trainable low-rank matrices. Similarly, prefix-tuning optimizes task-specific vectors appended to model inputs, mimicking "virtual tokens" to guide downstream tasks without altering core parameters [3]. While effective, these low-rank approximations inherently limit feature representation adaptability, impair convergence in complex tasks, and exhibit sensitivity to rank selection [4, 1, 5].
Recent quantum-inspired methodologies address these limitations through two complementary paradigms. The first leverages tensor-based adaptations, such as Quantum Tensor Adaptation (QuanTA) [6], which employs quantum circuit-inspired tensor decomposition for high-order parameter adjustments, and Tensor Product Attention (TPA) [7], which optimizes memory efficiency via contextual low-rank factorization. Parallel efforts, including CompactifAI [8], integrate tensor networks with singular value truncation for model compression. The second paradigm combines quantum neural network (QNN) [9] with classical architectures: Quantum-PEFT [10] achieves logarithmic parameter scaling through entangled unitary transformations, while Quantum Parameter Adaptation (QPA) [11] generates compact tuning parameters via hybrid quantum-classical mappings. These approaches integrate quantum-derived high-rank information into classical low-rank spaces, with Matrix Product Operator (MPO) representations [12] further enhancing robustness through localized entanglement regularization.
Despite these advances, existing quantum-inspired frameworks remain largely theoretical, lacking validation on physical quantum hardware. To bridge this gap, we propose the Quantum Weighted Tensor Hybrid Network (QWTHN), which reparameterizes pre-trained layers into quantum tensor hybrid architectures. By synergizing QNN for capturing complex transformations with tensor-based efficiency, our framework achieves superior parameter-efficient fine-tuning while surpassing conventional LoRA in performance. Our main contributions include:
First Implementation of Quantum Computing Inference on Real Machines: We propose a novel quantum large model fine-tuning algorithm, QWTHN, based on a hybrid quantum-classical neural network. The framework leverages the comprehensive learning capabilities of quantum neural networks and quantum tensor networks to achieve parameter-efficient fine-tuning (PEFT). For the first time, this approach enables inference technology on actual quantum computing hardware.
Significant Reduction in Trainable Parameters: The QWTHN framework reduces the number of trainable parameters by 76% compared to LoRA under equivalent conditions, achieving superior parameter efficiency while maintaining model performance.
Efficient Training: By significantly reducing the number of parameters, the model becomes easier to train. In certain scenarios, the convergence speed can be accelerated by 20% in terms of step size (i.e., the rate of loss value change per unit time), enabling the model to achieve superior fitting performance with fewer training steps. This not only mitigates the risk of overfitting but also holds potential for reducing training costs. Furthermore, the quantum tensor network emphasizes the local correlations within input signals, which helps prevent the training process from being trapped in suboptimal local minima.
Enhancing the Expressive Power of Current Low-Rank Fine-Tuning Networks: We demonstrate that QWTHN not only serves as a viable alternative to classical parameter-efficient fine-tuning methods such as LoRA and DoRA, but also exhibits superior performance across a broader spectrum of fine-tuning tasks. Specifically, on the CPsyCounD and R1-Distill-SFT datasets, QWTHN achieves up to a 15% reduction in training loss while significantly decreasing the number of trainable parameters. Moreover, it improves accuracy metrics by up to 20% on small-scale test sets through optimized parameter adaptation mechanisms.
This paper is structured as follows: Section I introduces the research background of large-scale language models, reviews recent advances in quantum-inspired methodologies for model adaptation, and outlines the significance of our work in addressing existing gaps. Section II critically examines classical fine-tuning techniques for large models and explores quantum-based approaches to integrating quantum principles into LLM adaptation frameworks. Section III details our experimental methodology, including the proposed model architecture, presents empirical results with comparative analyses, and discusses the implications of the findings for efficient model fine-tuning. Finally, Section IV summarizes the research contributions and proposes future directions for advancing quantum-augmented LLM technologies.
II Methods
QNN dynamically generate task-adapted weights through quantum entanglement and superposition, enabling finer-grained feature selection. Here, the MPO extracts foundational features, while the QNN focuses on generating optimized weights, achieving optimal computational resource allocation. Quantum circuits can explore a broader solution space [13], leveraging quantum state superposition to evaluate multiple weight combinations simultaneously, thereby circumventing the local optima limitations inherent in classical low-rank models.
II-A Quantum Tensor Network Based on MPO
To more efficiently represent the low-rank weight matrix W in LoRA, we employ a tensor decomposition method called the Matrix Product Operator (MPO) [12]. By reorganizing the elements of matrix W into a higher-dimensional tensor, we obtain a tensor with 2n indices, as show in Fig 1.
Figure 1: MPO decomposition of the weight matrix W. The local tensors $\mathbf{w}^{(k)}$ are represented by filled circles. The hollow circles denote the input and output indices $i_{l}$ and $j_{l}$, respectively, given the input index $i_{k}$ and the output index $j_{k}$, the slice $\mathbf{w}^{(k)}_{j_{k}i_{k}}$ of the local tensor $\mathbf{w}^{(k)}$ is a matrix.
Let the input space X and output space Y have dimensions $N_{x}$ and $N_{y}$, respectively. Through multilinear algebraic transformations, the original weight matrix $W\in\mathbb{R}^{N_{y}\times N_{x}}$ is mapped to a higher-order tensor:
$\mathcal{W}\in\mathbb{R}^{J_{1}\times\cdots\times J_{n}\times I_{1}\times\cdots\times I_{n}},$
the proposed method operates under the following dimensional constraints:
$\prod_{k=1}^{n}I_{k}=N_{x},\quad\prod_{k=1}^{n}J_{k}=N_{y}.$
Adopting a hierarchical index mapping strategy, we reshape the input vector $x\in\mathcal{X}$ into a multidimensional tensor $(i_{1},i_{2},\ldots,i_{n})$, with the output vector corresponding to $(j_{1},j_{2},\ldots,j_{n})$. We then decompose the LoRA weight matrix W into a product of local tensors ${w}^{(k)}$ via tensor factorization:
$\mathcal{W}_{j_{1}\cdots j_{n},i_{1}\cdots i_{n}}=\mathrm{Tr}\left[\prod_{k=1}^{n}\mathbf{w}^{(k)}_{j_{k}i_{k}}\right],$ (1)
where in the local tensor $\mathbf{w}^{(k)}\in\mathbb{R}^{D_{k-1}\times J_{k}\times I_{k}\times D_{k}}$ satisfies the following conditions:
$\mathbf{w}^{(k)}_{j_{k}i_{k}}:D_{k-1}\rightarrow D_{k}.$ (2)
Bond dimension parameter $D=\max\{D_{k}\}$ governs the model’s expressive capacity, with its value positively correlated to the quantum entanglement entropy. By establishing a controllable balancing mechanism between model complexity and expressive power, this approach provides a novel technical pathway for lightweight design of large-scale neural networks. For parameter optimization, the total number of trainable parameters satisfies:
$N_{\text{MPO}}=\sum_{k=1}^{n}I_{k}J_{k}D_{k-1}D_{k},$ (3)
when a uniform bond dimension $D_{k}\equiv D$ is adopted, the total parameter count simplifies to:
$N_{\text{MPO}}=D(I_{1}J_{1}+I_{n}J_{n})+D^{2}\sum_{k=2}^{n-1}I_{k}J_{k}.$ (4)
Compared to the $N_{x}\times N_{y}$ parameters of traditional fully connected layers, our method achieves exponential compression when $n\geq 3$. The initialization strategy employs an improved Kaiming uniform distribution[14], this method of initialization effectively avoids parameter redundancy during the early stages of training while maintaining gradient stability.
II-B QWTHN
The core of quantum neural networks lies in their ability to leverage the superposition and entanglement of quantum states, enabling efficient representation of high-dimensional features in Hilbert space. Assuming input data $x\in\mathbb{R}^{n}$ is encoded as a quantum state $|\psi(x)\rangle$, the system evolves into a new quantum state $|\phi(x)\rangle$ after a series of quantum gate operations. By measuring $|\phi(x)\rangle$, a classical output $y\in\mathbb{R}^{m}$ can be obtained. This process can be described by the following formula:
$y=\mathcal{M}(|\phi(x)\rangle),$ (5)
where $\mathcal{M}(\cdot)$ denotes the measurement operation. Although $y$ is classical data, its generation process involves nonlinear transformations of quantum states, enabling the extraction of features that are difficult to capture with classical methods. While classical neural networks can generate non-harmonic functions, QNN excel at fitting truncated Fourier series [15]. Therefore, the quantum-weighted hybrid network not only learns smooth features but also fills non-harmonic gaps through classical neural networks, significantly reducing training error and improving generalization.
To address the limited expressive power of classical neural networks (including MPO neural networks) in low-rank spaces, we perform element-wise multiplication of the classical output $W_{q}$ from the QNN with the output Y from the classical neural network:
$\tilde{O}=W_{q}\odot Y,$ (6)
where $\odot$ denotes the element-wise multiplication operation. The updated output $\tilde{O}$ retains the low-rank structure while incorporating high-dimensional features extracted by the QNN, thereby enhancing the model’s ability to capture complex nonlinear relationships. Through the element-wise multiplication, the QNN output $W_{q}$ provides additional nonlinear features to $Y$. The elements in $Y$ contain high-order feature interaction information related to the input data [16, 17, 18, 19], which is often neglected in low-rank approximations. Therefore, the expressive power of the updated output $\tilde{O}$ is theoretically improved.
Based on the aforementioned MPO and quantum-weighted hybrid network, we construct the final QWTHN, as shown in Fig 2.
Figure 2: QWTHN. The QWTHN leverages the MPO neural network to reduce the parameter count of LoRA. The input vector of LoRA is reshaped and fed into $\mathrm{MPO_{A}}$, where the last dimension of the output is output mpo $\mathrm{output}_{\mathrm{mpo_{A}}}$=128. This output is then transformed by a classical Multilayer Perceptron ($\mathrm{MLP_{A}}$) into output $\mathrm{output}_{\mathrm{MLP_{A}}}$=4, which corresponds to the number of qubits in the subsequent QNN.The output $\mathrm{output}_{\mathrm{MLP_{A}}}$ is input into the QNN, where it undergoes RY angle encoding and variational quantum state transformations. By exploiting the superposition and entanglement of quantum states, the QNN efficiently represents high-dimensional features in Hilbert space. Finally, the expectation values of PauliZ measurements on each qubit yield an output vector $\mathrm{output}_{\mathrm{QNN}}$ of the same size as $\mathrm{output}_{\mathrm{MLP_{A}}}$. The input to the next layer $\mathrm{MLP_{B}}$ is computed as: $\mathrm{input}_{\mathrm{MLP_{B}}}=\mathrm{output}_{\mathrm{QNN}}\odot\mathrm{output}_{\mathrm{MLP_{A}}}+\mathrm{output}_{\mathrm{MLP_{A}}}$, where $\mathrm{output}_{\mathrm{QNN}}\odot\mathrm{output}_{\mathrm{MLP_{A}}}$ corresponds to Equation (6). The $\mathrm{MLP_{B}}$ then linearly transforms input $\mathrm{output}_{\mathrm{MLP_{B}}}$ into $\mathrm{output}_{\mathrm{MLP_{B}}}$=128, which is fed into $\mathrm{MPO_{B}}$ to produce the final output of the LoRA layer based on QWTHN. All the above dimensions can be set as hyperparameters.
We employ RY rotation angle encoding to reduce the number of encoding layers, thereby enhancing the training efficiency of the QNN and improving the scalability for real quantum hardware execution. By utilizing parameterized two-qubit gates CRZ (Controlled-Rotation-Z), we can adjust parameters to enhance the exploration of the state space. Our design adopts a block structure, where each block consists of continuous nearest-neighbor interactions and one non-local interaction. This modular approach supports performance enhancement through stacking multiple layers, yet achieves near-saturation of expressive power with fewer layers, thus reducing optimization complexity. This design is particularly well-suited for current quantum computing hardware.
III Experiment
III-A Datasets
The datasets used in this study are publicly available, specifically the CPsyCoun[20] Chinese dataset, the R1-Distill-SFT[21] English dataset and the Chinese-DeepSeek-R1-Distill-data-110k[22].
The CPsyCoun Chinese dataset is based on the open-source version available on the HuggingFace platform, released by the CAS-SIAT-XHai team. It is a psychological counseling dialogue dataset containing real-world Chinese dialogue records from counseling scenarios, covering nine common psychological issues such as depression, anxiety, and interpersonal relationships. The dataset spans seven classic theoretical schools in the field of psychological counseling, representing different theoretical frameworks in counseling practice, and is considered a representative branch of technical methods. The data has been anonymized and annotated with multi-turn dialogue roles between counselors and clients, comprising approximately 16,000 dialogue samples. Some data includes fine-grained labels such as counseling strategies and types of emotional support. This dataset is suitable for research on psychological counseling dialogue generation, sentiment analysis, and intervention strategy modeling, providing a benchmark for exploring the application of natural language processing technologies in the mental health domain.
The R1-Distill-SFT dataset is one of the open-source datasets provided by the ModelScope community. It is a structured question bank aimed at K-12 mathematics education, covering areas such as arithmetic, algebra, geometry, and comparison, along with step-by-step solutions. Each data sample includes the question stem, blanks to be filled, multilingual labels (e.g., English and Chinese cn_k12), standardized answers (annotated using LaTeX mathematical symbols), and verification status. The dataset is suitable for model training (e.g., Supervised Fine-Tuning, SFT) and automated evaluation in educational scenarios. It has been filtered based on quality metrics, covering topics ranging from basic operations to medium complexity, and supports research on personalized learning and reasoning ability assessment. The R1-Distill-SFT dataset focuses on supporting tasks such as Knowledge Distillation and Supervised Fine-Tuning (SFT) for models.
The Chinese-DeepSeek-R1-Distill-data-110k is one of the Chinese datasets provided by the ModelScope community. It is a Chinese instruction fine-tuning dataset generated by distilling DeepSeek-R1, containing approximately 110,000 high-quality multi-turn dialogue samples. This dataset is designed to support the fine-tuning and optimization of the DeepSeek-R1 series of models. It covers a wide range of domains, including general dialogue, knowledge-based question answering, code generation, and mathematical reasoning. The dataset is generated using distillation techniques from large language models and undergoes rigorous cleaning and deduplication processes. The data is structured into input-output pairs with contextual annotations, ensuring high diversity in instructions, strong logical coherence, and broad task coverage. It is particularly suitable for training dialogue models capable of understanding and executing complex instructions.
III-B Evaluation Metrics
In this experiment, Cross-Entropy Loss[23] is adopted as the evaluation metric for the model. Cross-entropy measures the difference between the predicted probability distribution of the model and the true label distribution, effectively reflecting the optimization level of classification tasks. A smaller value indicates that the model’s predictions are closer to the true distribution. For a dataset containing N samples, the cross-entropy loss is calculated as:
$\mathcal{L}=-\frac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{C}y_{ik}\log(p_{ik}),$
where C represents the total number of classes in the classification task, $y_{ik}\in\{0,1\}$ is the true label of sample i for class k (in one-hot encoding form), and $p_{ik}\in(0,1)$ denotes the predicted probability that sample i belongs to class k. This loss function averages the prediction errors across all samples through the double summation operation, effectively penalizing misclassified predictions while ensuring comparability of losses across different batches of data.
For text generation evaluation, we utilize BLEU-4[24] and ROUGE[25] scores. BLEU-4 calculates 4-gram precision with a brevity penalty BP to penalize short outputs:
[…]
$\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases},$
$\text{Precision}_{n}=\frac{\sum\text{Count}_{\text{clip}}(\text{n-gram})}{\sum\text{Count}_{\text{generated}}(\text{n-gram})},$
  • BLEU-4: Bilingual Evaluation Understudy score for 4-gram precision,
  • BP: Brevity Penalty to penalize short translations,
  • $w_{n}$: Uniform weights with $w_{n}=\frac{1}{4}$,
  • $\text{Precision}_{n}$: Modified n-gram precision for $n=1$ to $4$,
  • $c$: Total length of candidate translations,
  • $r$: Effective reference corpus length,
[…]
The recall metric ROUGE-N quantifies the proportion of n-grams in the reference(s) text that are successfully reproduced in the generated text. Formally, it is computed as:
[…]
  • R: the set of reference text(s),
[…]
The ROUGE-L metric evaluates text generation quality by measuring the alignment between a generated text and reference text(s) using the Longest Common Subsequence (LCS). Unlike n-gram-based metrics, ROUGE-L emphasizes semantic coherence and structural continuity by identifying the longest sequence of tokens that appear in both texts in the same order, though not necessarily contiguously. The metric is computed as a balanced F1-score derived from LCS-based recall ($R_{\text{LCS}}$) and precision ($P_{\text{LCS}}$):
[…]
$R_{\text{LCS}}=\frac{\text{Length of LCS(Ref,Gen)}}{\text{Total tokens in reference text(s)}},$
$P_{\text{LCS}}=\frac{\text{Length of LCS(Ref,Gen)}}{\text{Total tokens in generated text}},$
  • $R_{\text{LCS}}$: Recall of the Longest Common Subsequence,
  • $P_{\text{LCS}}$: Precision of the Longest Common Subsequence,
  • $\beta$: Controls the relative weight of recall ($R_{\text{LCS}}$) and precision ($P_{\text{LCS}}$). Typically, $\beta$=1 to prioritize harmonic mean (F1).
III-C Based Model
The baseline model for comparison is the LoRA algorithm with a rank of 4, fine-tuned based on the DeepSeek-R1-Distill-Qwen-7B (https://modelscope.cn/models/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B) and Qwen2-7B-Instruct models.
III-D Results
This section primarily aims to verify whether the efficient fine-tuning algorithm described for the QWTHN can improve the performance of fine-tuned models while significantly reducing the number of parameters. This evaluation is based on the assumption that the QWTHN can enhance the expressive power of classical neural networks in low-rank spaces. The experiments were conducted using PyTorch[26] and PyVQNet[27, 28] with quantum circuit simulations. In this stage, noise effects in the quantum system are ignored, and the Cross-Entropy Loss is adopted as the optimization objective for classification tasks. As shown in Fig 3, the convergence curve of QWTHN exhibits a steeper descent rate during early training stages and achieves a lower stable loss value in training phase.
Figure 3: Training loss comparison of QWTHN, QuanTA, and LoRA across varying sample sizes (N) and datasets. Subplots (a–c) illustrate results on the CPsyCounD dataset with N={300,600,1000}, while subplots (d–f) show corresponding analyses on the R1-Distill-SFT dataset. QWTHN consistently achieves lower training losses compared to baseline methods, demonstrating enhanced convergence efficiency under diverse experimental conditions.
The datasets are extracted from CPsyCounD and R1-Distill-SFT with 300, 600, and 1000 samples, respectively. The maximum sequence length max_seq_len for both training and prediction is set to 1024, with 10% of the data used as the validation set and the remaining 90% as the training set. For the Qwen2-7B-Instruct test set, the last 50 samples from CPsyCounD (with no overlap with the previous data) are used as the test set.
Since DeepSeek-R1-Distill-Qwen-7B excels in code and mathematical capabilities, using metrics like BLEU-4 to compare the performance of QWTHN LoRA and MLP LoRA is not appropriate. Instead, we evaluate the test set performance using the following metrics:
Strict Accuracy (SA): Defined as whether the output contains the correct answer in the answer segment. Accuracy: Unlike SA, this metric considers the answer correct if the correct answer appears anywhere in the reasoning chain. Chain-of-Thought Completeness (CTC): Evaluates whether the response includes a complete reasoning chain, assessing the model’s ability to apply reasoning chains within limited lengths. Answer Completeness (AC): Measures the completeness of the answer itself.
As demonstrated in Table I, the proposed QWTHN achieves significantly higher training efficiency than classical LoRA under identical fine-tuning conditions for LLMs. When evaluated on chain-of-thought (CoT) reasoning tasks, QWTHN exhibits statistically significant improvements in domain-specific metrics—including Contextual Task Consistency (CTC), Answer Coherence (AC), Accuracy, and Strict Accuracy—while reducing trainable parameters by over 70% compared to LoRA.
Further validation across benchmark datasets (CPsyCounD, R1-Distill-SFT, and Chinese-DeepSeek-R1-Distill-data-110k) confirms the robustness of QWTHN. As summarized in Table II, the framework demonstrates consistent gains in text generation quality, with absolute increases of 1–3 points in BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L metrics. These results underscore QWTHN’s ability to reconcile parameter efficiency with enhanced model performance, establishing it as a viable alternative to conventional low-rank adaptation paradigms.
Table I: Cross-Architecture Performance Comparison by Dataset Scale
CPsyCounD (Training/Validation Loss)
Model Small (300) Medium (600) Full (1000)
Train Val Train Val Train Val
MLP 1.805 1.606 1.657 1.514 1.577 1.447
quanTA 1.647 1.512 1.565 1.478 1.520 1.437
QWTHN 1.529 1.475 1.512 1.464 1.497 1.464
R1-Distill-SFT (Training/Validation Loss)
Model Small (300) Medium (600) Full (1000)
Train Val Train Val Train Val
MLP 0.179 0.182 0.172 0.158 0.169 0.160
quanTA 0.167 0.173 0.164 0.159 0.163 0.159
QWTHN 0.155 0.174 0.158 0.166 0.162 0.164
Note. Parameter counts: MLP = 1.26M (100%), quanTA = 1.37M (108.7%), QWTHN = 0.30M (23.7%).
Table II: Performance Metrics for CPsyCounD, R1-Distill-SFT and Chinese-DeepSeek-R1-Distill-data-110k
CPsyCounD
Model BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
MLP 13.459 34.540 13.109 30.627
quanTA 8.650 29.463 9.780 25.539
QWTHN 14.593 37.345 14.023 32.32
R1-Distill-SFT
Model BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
MLP 59.635 64.195 42.627 43.326
quanTA 59.096 63.512 41.992 42.227
QWTHN 61.615 66.087 45.294 45.915
Chinese-DeepSeek-R1-Distill-data-110k
Model CTC AC Accuracy Strict Accuracy
MLP 78% 68% 82% 68%
quanTA 84% 74% 82% 68%
QWTHN 96% 94% 86% 82%
III-E Quantum Running
We have constructed a variational quantum circuit module incorporating the Origin Quantum Cloud (QCloudService) in our model architecture. This module utilizes the quantum chip backend codenamed "origin_wukong" to transform user-defined variational quantum circuits (including input data and trainable parameters) into quantum circuit intermediate representations (originIR) compatible with the QPanda framework[29]. Given that Origin Quantum chips exclusively support Z-axis observable measurements, basis transformation operations are implemented to convert user-defined non-Z observables into equivalent Z-axis measurements. Specifically, each Pauli operator measurement generates a distinct quantum circuit.
Furthermore, to accommodate multi-dimensional input data (where input dimensions may exceed unity), each data dimension is entangled with trainable parameters in the variational quantum circuit to form a complete quantum architecture. Consequently, the total number of quantum circuits executed per batch on the quantum hardware is governed by the relationship:
$\mathrm{T=B\times Q},$
where B denotes the batch size and Q represents the number of qubits in the system.
To comply with quantum hardware batch processing constraints, we leverage the backend’s asynchronous batch measurement interface run to partition quantum circuits into groups for sequential submission to QCloudService. As run operates asynchronously, we continuously monitor task status through the returned QCloudResult object until all measurement outcomes are retrieved.
Finally, measurement results are reformatted into machine learning-compatible tensors through concatenation and reshaping operations. These processed outputs can be directly utilized by subsequent neural network modules, establishing seamless integration between quantum computations and classical machine learning pipelines.
III-F Discussion
The parameter count of the Quantum Tensor Hybrid Network (QWTHN), which combines quantum tensor networks and quantum neural networks, is lower than that of classical LoRA, while the training loss is significantly reduced. This leads to improved parameter efficiency. The Matrix Product Operator (MPO) decomposition introduces the concept of rank, effectively decomposing high-dimensional linear transformations into a series of low-dimensional matrix products. This decomposition significantly reduces the number of model parameters, especially when handling high-dimensional inputs and outputs.
Traditional fully connected layers flatten the input into vectors, losing the dimensional structure of the original data. In contrast, MPO neural networks preserve this structural information to some extent because they operate on the dimensions of the original tensor. This structured approach may enable MPO to learn and represent multidimensional relationships and patterns in the data more effectively, such as correlations between dimensions and hierarchical structures.
By leveraging the classical output of the QNN to perform element-wise multiplication with the output of the classical network, the expressive power of classical neural networks in low-rank spaces is enhanced. Quantum neural networks process input data using quantum properties (e.g., superposition, entanglement, and parallel computation), potentially generating high-dimensional or nonlinear features. Even though the final output of the QNN is classical data (e.g., probability distributions after measurement), the underlying quantum state evolution process can still extract information that is difficult to obtain using classical methods.
Currently, the scale of dataset experiments is limited by the size of quantum circuits. Although it can be seen that training loss on a larger dataset of 1000 samples still performs better. Future work will need to focus on developing fine-tuning strategies for larger datasets and even exploring large-scale quantum pre-training schemes. The current integration of MPO (Matrix Product Operator) and VQC (Variational Quantum Circuit) involves hard connections between input and output interactions. However, theoretically, all weight matrices could be replaced with variational quantum circuits, which would enable the model to capture higher degrees of correlations far beyond what classical weight matrices can represent. Weight matrix W can be decomposed into an MPO with a given bond dimension, and this MPO can be further disentangled using quantum circuits, with the remaining operators also expressed in MPO format [30]. This is one feasible technical direction, and we will explore more potential implementations of large-scale quantum models.
IV Conclusion
This paper implements a quantum-weighted tensor hybrid fine-tuning algorithm for large models, leveraging quantum computing principles to significantly reduce parameter redundancy and enhance parameter efficiency. It validates the potential of quantum computing to alleviate the computational bottleneck of large models, achieving a novel integration of quantum computing and large-scale models. At the current stage of quantum computer development, this work not only demonstrates the efficient adaptation of quantum resources to billion-parameter models for lightweight optimization but also validates a practical pathway for quantum hardware driven by large model tasks. This lays the first engineerable technical foundation for future quantum-enhanced Artificial General Intelligence (AGI) systems.

### v2

@@
 Abstract
-Low-Rank Adaptation (LoRA) enables efficient fine-tuning of pre-trained language models via low-rank matrix approximation, which is effective in many scenarios.
-However, its low-rank representation capacity is constrained in complex tasks or high-rank dependency settings, potentially limiting model adaptability.
-Addressing the expressive bottleneck of classical low-rank approximation in fine-tuning large language models, this paper proposes a parameter-efficient fine-tuning method based on a Quantum Weighted Tensor Hybrid Network (QWTHN), which leverages Quantum Neural Network (QNN).
-The study investigates quantum-classical hybrid parameter-efficient fine-tuning in low-rank spaces.
-QWTHN decomposes pre-trained weights into quantum neural network and tensor network representations, utilizing quantum state superposition and other methods to break through classical rank limitations.
-Experiments show that the proposed quantum fine-tuning technique for large models approaches or even surpasses the parameter efficiency of LoRA.
-On the CPsyCounD and R1-Distill-SFT datasets, QWTHN, compared to classical LoRA, reduces training loss by up to 15% while using 76% fewer parameters, and achieves an 8.4% performance improvement on the CPsyCounD test set.
-This research not only realizes lightweight and efficient adaptation of quantum resources to billion-parameter models but also validates the practical path of quantum hardware driven by large model tasks, laying the first engineering-ready technical foundation for future quantum-enhanced AGI systems.
+Low-Rank Adaptation (LoRA) enables efficient fine-tuning of pre-trained language models through low-rank matrix approximation, achieving effectiveness in many scenarios.
+However, its representation capacity is constrained in complex tasks or high-rank dependency settings, potentially limiting model adaptability.
+To overcome the expressive bottleneck in classical low-rank approximation for fine-tuning large language models (LLMs), we propose Quantum Tensor Hybrid Adaptation (QTHA), a parameter-efficient fine-tuning method that integrates a quantum neural network (QNN) with a tensor network.
+QTHA explores quantum tensor hybrid fine-tuning within low-rank spaces by decomposing pre-trained weights into quantum neural network and tensor network representations, leveraging quantum state superposition to overcome classical rank limitations.
+Experiments demonstrate that QTHA achieves performance comparable to or surpassing LoRA in parameter-efficient fine-tuning.
+Compared to LoRA, QTHA reduces trainable parameters by 76% while reducing training loss by up to 17% and improving test set performance by up to 17% within the same training steps.
+This research not only enables lightweight adaptation of quantum resources to the billion-parameter models but also validates the feasibility of quantum hardware optimization driven by LLM tasks.
+It establishes the first engineering-ready foundation for future quantum-enhanced Artificial General Intelligence (AGI) systems.
 Index Terms:
-Quantum Computing; LLM; Quantum LLM; Fine Tuning
+Quantum Computing; LLM; Quantum LLM; Fine-Tuning
 I Introduction
 The rapid advancement of large language models (LLMs) has driven innovations in parameter-efficient fine-tuning (PEFT) to reduce computational overhead while preserving performance.
-Classical methods such as Low-Rank Adaptation (LoRA) [1] and DoRA [2] operate under the assumption that weight updates during fine-tuning reside in low-rank subspaces, enabling efficient adaptation through trainable low-rank matrices.
-Similarly, prefix-tuning optimizes task-specific vectors appended to model inputs, mimicking "virtual tokens" to guide downstream tasks without altering core parameters [3].
+Classical methods such as Low-Rank Adaptation (LoRA) [1] and Weighted-Decomposed Low-Rank Adaptation (DoRA) [2] assume that weight updates during fine-tuning lie within low-rank subspaces, enabling efficient adaptation via trainable low-rank matrices.
+Similarly, prefix-tuning optimizes task-specific vectors appended to model inputs, mimicking “virtual tokens” to guide downstream tasks without altering core parameters [3].
 While effective, these low-rank approximations inherently limit feature representation adaptability, impair convergence in complex tasks, and exhibit sensitivity to rank selection [1, 2, 4].
@@
 Despite these advances, existing quantum-inspired frameworks remain largely theoretical, lacking validation on physical quantum hardware.
-To bridge this gap, we propose the Quantum Weighted Tensor Hybrid Network (QWTHN), which reparameterizes pre-trained layers into quantum tensor hybrid architectures.
+To bridge this gap, we propose the Quantum Tensor Hybrid Adaptation (QTHA), which reparameterizes pre-trained layers into quantum tensor hybrid architectures.
 By synergizing QNN for capturing complex transformations with tensor-based efficiency, our framework achieves superior parameter-efficient fine-tuning while surpassing conventional LoRA in performance.
 Our main contributions include:
-First Implementation of Quantum Computing Inference on Real Machines: We propose a novel quantum large model fine-tuning algorithm, QWTHN, based on a hybrid quantum-classical neural network.
-The framework leverages the comprehensive learning capabilities of quantum neural networks and quantum tensor networks to achieve parameter-efficient fine-tuning (PEFT).
-For the first time, this approach enables inference technology on actual quantum computing hardware.
-Significant Reduction in Trainable Parameters: The QWTHN framework reduces the number of trainable parameters by 76% compared to LoRA under equivalent conditions, achieving superior parameter efficiency while maintaining model performance.
-Efficient Training: By significantly reducing the number of parameters, the model becomes easier to train.
-In certain scenarios, the convergence speed can be accelerated by 20% in terms of step size (i.e., the rate of loss value change per unit time), enabling the model to achieve superior fitting performance with fewer training steps.
-This not only mitigates the risk of overfitting but also holds potential for reducing training costs.
-Furthermore, the quantum tensor network emphasizes the local correlations within input signals, which helps prevent the training process from being trapped in suboptimal local minima.
-Enhancing the Expressive Power of Current Low-Rank Fine-Tuning Networks: We demonstrate that QWTHN not only serves as a viable alternative to classical parameter-efficient fine-tuning methods such as LoRA and DoRA, but also exhibits superior performance across a broader spectrum of fine-tuning tasks.
-Specifically, on the CPsyCounD and R1-Distill-SFT datasets, QWTHN achieves up to a 15% reduction in training loss while significantly decreasing the number of trainable parameters.
-Moreover, it improves accuracy metrics by up to 20% on small-scale test sets through optimized parameter adaptation mechanisms.
-This paper is structured as follows: Section I introduces the research background of large-scale language models, reviews recent advances in quantum-inspired methodologies for model adaptation, and outlines the significance of our work in addressing existing gaps.
-Section II critically examines classical fine-tuning techniques for large models and explores quantum-based approaches to integrating quantum principles into LLM adaptation frameworks.
-Section III details our experimental methodology, including the proposed model architecture, presents empirical results with comparative analyses, and discusses the implications of the findings for efficient model fine-tuning.
-Finally, Section IV summarizes the research contributions and proposes future directions for advancing quantum-augmented LLM technologies.
+First implementation of quantum computing inference for LLM on quantum hardware: We introduce QTHA, a novel quantum-enhanced fine-tuning algorithm for LLM, based on a hybrid quantum-classical neural network architecture.
+This framework synergistically combines the expressive power of QNN with the efficiency of tensor networks to achieve PEFT.
+Notably, QTHA represents the first practical implementation of inference technology on quantum hardware, bridging the gap between theoretical quantum machine learning and deployable solutions.
+Significant reduction in trainable parameters: The QTHA method demonstrates remarkable parameter efficiency, reducing trainable parameters by 76% compared to LoRA under identical conditions while preserving model performance.
+This substantial parameter reduction significantly enhances model trainability.
+Notably, in specific scenarios, QTHA accelerates convergence by 20% in terms of step size (measured by the rate of loss value change per unit time), enabling the model to achieve optimal fitting performance with fewer training iterations.
+This dual advantage not only reduces overfitting risks but also shows promising potential for lowering computational costs.
+Enhancing the performance of LLM fine-tuning: We demonstrate that QTHA not only serves as a viable alternative to classical parameter-efficient fine-tuning methods such as LoRA and DoRA, but also exhibits superior performance across a broader spectrum of fine-tuning tasks.
+Specifically, QTHA achieves up to a 17% reduction in training loss while significantly decreasing the number of trainable parameters.
+Moreover, it improves accuracy metrics by up to 17% on small-scale test sets through optimized parameter adaptation mechanisms.
+This study employs a systematic framework to advance quantum-enhanced language model research: Section I establishes the innovative potential of this work in addressing domain-critical challenges by tracing the evolutionary trajectory of LLMs and synthesizing advancements in quantum-inspired adaptation methodologies.
+Section II deconstructs conventional fine-tuning paradigms while proposing a groundbreaking quantum-integrated strategy that architecturally embeds quantum computational principles into LLM adaptation frameworks.
+Section III designs an innovative hybrid experimental protocol, demonstrating statistically significant benchmarking results through comprehensive comparative analyses to elucidate the mechanism of resource-efficient quantum-enhanced fine-tuning.
+The concluding Section IV crystallizes dual-impact theoretical-practical innovations in quantum-enhanced fine-tuning of LLMs, charting transformative pathways for next-generation LLM ecosystems.
 II Methods
-QNN dynamically generate task-adapted weights through quantum entanglement and superposition, enabling finer-grained feature selection.
-Here, the MPO extracts foundational features, while the QNN focuses on generating optimized weights, achieving optimal computational resource allocation.
-Quantum circuits can explore a broader solution space [13], leveraging quantum state superposition to evaluate multiple weight combinations simultaneously, thereby circumventing the local optima limitations inherent in classical low-rank models.
+The QNN leverages the dynamic properties of quantum entanglement and superposition to achieve highly nonlinear feature modeling capabilities.
+In this framework, the MPO is responsible for efficiently extracting abrupt features, while the QNN focuses on effectively learning periodic features[12, 11].
+Through the linear combination of the two, this approach aims to overcome the limitations of classical linear layers in feature learning, thereby achieving optimal allocation of computational resources.
+Furthermore, quantum circuits can explore a broader solution space[13], circumventing the inherent limitations of local optima of classical low-rank models.
 II-A Quantum Tensor Network Based on MPO
-To more efficiently represent the low-rank weight matrix W in LoRA, we employ a tensor decomposition method called the Matrix Product Operator (MPO) [12].
-By reorganizing the elements of matrix W into a higher-dimensional tensor, we obtain a tensor with 2n indices, as show in Fig 1.
-Figure 1: MPO decomposition of the weight matrix W.
-The local tensors $\mathbf{w}^{(k)}$ are represented by filled circles.
-The hollow circles denote the input and output indices $i_{l}$ and $j_{l}$, respectively, given the input index $i_{k}$ and the output index $j_{k}$, the slice $\mathbf{w}^{(k)}_{j_{k}i_{k}}$ of the local tensor $\mathbf{w}^{(k)}$ is a matrix.
-Let the input space X and output space Y have dimensions $N_{x}$ and $N_{y}$, respectively.
-Through multilinear algebraic transformations, the original weight matrix $W\in\mathbb{R}^{N_{y}\times N_{x}}$ is mapped to a higher-order tensor:
-$\mathcal{W}\in\mathbb{R}^{J_{1}\times\cdots\times J_{n}\times I_{1}\times\cdots\times I_{n}},$
-the proposed method operates under the following dimensional constraints:
-$\prod_{k=1}^{n}I_{k}=N_{x},\quad\prod_{k=1}^{n}J_{k}=N_{y}.$
+To efficiently represent the low-rank weight matrix $\mathbf{W}$ in LoRA, we employ a tensor decomposition method called the Matrix Product Operator (MPO) [11].
+By reorganizing the elements of matrix $\mathbf{W}$ into a higher-dimensional tensor, we obtain a tensor with $2n$ indices.
+Specifically, let the input space $\mathcal{X}$ and output space $\mathcal{Y}$ have dimensions $N_{x}$ and $N_{y}$, respectively.
+Through multilinear algebraic transformations, the original weight matrix $\mathbf{W}\in\mathbb{R}^{N_{y}\times N_{x}}$ is mapped to a higher-order tensor:
+$\mathcal{W}\in\mathbb{R}^{J_{1}\times\cdots\times J_{n}\times I_{1}\times\cdots\times I_{n}},$ (1)
+where
+$\prod_{k=1}^{n}I_{k}=N_{x},\quad\prod_{k=1}^{n}J_{k}=N_{y}.$ (2)
 Adopting a hierarchical index mapping strategy, we reshape the input vector $x\in\mathcal{X}$ into a multidimensional tensor $(i_{1},i_{2},\ldots,i_{n})$, with the output vector corresponding to $(j_{1},j_{2},\ldots,j_{n})$.
-We then decompose the LoRA weight matrix W into a product of local tensors ${w}^{(k)}$ via tensor factorization:
-$\mathcal{W}_{j_{1}\cdots j_{n},i_{1}\cdots i_{n}}=\mathrm{Tr}\left[\prod_{k=1}^{n}\mathbf{w}^{(k)}_{j_{k}i_{k}}\right],$ (1)
+We then decompose the weight matrix $\mathbf{W}$ into a product of local tensors ${w}^{(k)}$ via tensor factorization:
+$\mathcal{W}_{j_{1}\cdots j_{n},i_{1}\cdots i_{n}}=\mathrm{Tr}\left[\mathbf{w}^{(1)}_{j_{1}i_{1}}\mathbf{w}^{(2)}_{j_{2}i_{2}}\cdots\mathbf{w}^{(n)}_{j_{n}i_{n}}\right],$ (3)
 where in the local tensor $\mathbf{w}^{(k)}\in\mathbb{R}^{D_{k-1}\times J_{k}\times I_{k}\times D_{k}}$ satisfies the following conditions:
-$\mathbf{w}^{(k)}_{j_{k}i_{k}}:D_{k-1}\rightarrow D_{k}.$ (2)
-Bond dimension parameter $D=\max\{D_{k}\}$ governs the model’s expressive capacity, with its value positively correlated to the quantum entanglement entropy.
+$\mathbf{w}^{(k)}_{j_{k}i_{k}}\in\mathbb{R}^{D_{k-1}\times D_{k}}.$ (4)
+Here, bond dimension parameter $D=\max\{D_{k}\}$ governs the model’s expressive capacity, with its value positively correlated to the quantum entanglement entropy.
 By establishing a controllable balancing mechanism between model complexity and expressive power, this approach provides a novel technical pathway for lightweight design of large-scale neural networks.
@@
 $N_{\text{MPO}}=\sum_{k=1}^{n}I_{k}J_{k}D_{k-1}D_{k},$ (5)
-when a uniform bond dimension $D_{k}\equiv D$ is adopted, the total parameter count simplifies to:
+when a uniform bond dimension $D_{k}\equiv D$ is adopted, and the total parameter count simplifies to:
 $N_{\text{MPO}}=D(I_{1}J_{1}+I_{n}J_{n})+D^{2}\sum_{k=2}^{n-1}I_{k}J_{k}.$ (6)
@@
 The initialization strategy employs an improved Kaiming uniform distribution [14], this method of initialization effectively avoids parameter redundancy during the early stages of training while maintaining gradient stability.
-II-B QWTHN
+II-B Quantum Tensor Hybrid Adaptation
 The core of quantum neural networks lies in their ability to leverage the superposition and entanglement of quantum states, enabling efficient representation of high-dimensional features in Hilbert space.
-Assuming input data $x\in\mathbb{R}^{n}$ is encoded as a quantum state $|\psi(x)\rangle$, the system evolves into a new quantum state $|\phi(x)\rangle$ after a series of quantum gate operations.
-By measuring $|\phi(x)\rangle$, a classical output $y\in\mathbb{R}^{m}$ can be obtained.
-This process can be described by the following formula:
-$y=\mathcal{M}(|\phi(x)\rangle),$ (5)
-where $\mathcal{M}(\cdot)$ denotes the measurement operation.
+We designed a highly expressive QNN, in which an input vector $x\in\mathbb{R}^{n}$ is assigned to a quantum state $|\psi(x)\rangle$ via angle embedding using $R_{Y}$ gates:
+$|\psi(x)\rangle=\bigotimes_{i=1}^{q}R_{Y}(x_{i})|0\rangle,$ (7)
+where $q$ denotes the number of qubits.
+The quantum state evolves through a parameterized unitary transformation $U(\theta)$:
+$|\phi(x)\rangle=U(\theta)|\psi(x)\rangle=\prod_{l=1}^{L}\left(\bigotimes_{i=1}^{q}R_{Y}(\theta_{l,i})\cdot\text{CRZ}(\theta_{l,ij})\right),$ (8)
+where $L$ is the number of layers. CRZ represents controlled-Z rotation gates, which introduces entanglement between qubits.
+Each qubit is measured through Pauli-Z observables, yielding expectation values:
+$y_{i}=\langle\phi(x)|Z_{i}|\phi(x)\rangle\in[-1,1],$ (9)
+with the final output $O_{q}=[y_{1},\ldots,y_{m}]$ being a classical vector.
+By leveraging quantum state superposition and entanglement, QNN generate nonlinear features in high-dimensional Hilbert space.
 Although $y$ is classical data, its generation process involves nonlinear transformations of quantum states, enabling the extraction of features that are difficult to capture with classical methods.
-While classical neural networks can generate non-harmonic functions, QNN excel at fitting truncated Fourier series [15].
-Therefore, the quantum-weighted hybrid network not only learns smooth features but also fills non-harmonic gaps through classical neural networks, significantly reducing training error and improving generalization.
-To address the limited expressive power of classical neural networks (including MPO neural networks) in low-rank spaces, we perform element-wise multiplication of the classical output $W_{q}$ from the QNN with the output Y from the classical neural network:
-$\tilde{O}=W_{q}\odot Y,$ (6)
-where $\odot$ denotes the element-wise multiplication operation.
-The updated output $\tilde{O}$ retains the low-rank structure while incorporating high-dimensional features extracted by the QNN, thereby enhancing the model’s ability to capture complex nonlinear relationships.
-Through the element-wise multiplication, the QNN output $W_{q}$ provides additional nonlinear features to $Y$.
-The elements in $Y$ contain high-order feature interaction information related to the input data [16, 17, 18, 19], which is often neglected in low-rank approximations.
-Therefore, the expressive power of the updated output $\tilde{O}$ is theoretically improved.
-Based on the aforementioned MPO and quantum-weighted hybrid network, we construct the final QWTHN, as shown in Fig 2.
-Figure 2: QWTHN.
-The QWTHN leverages the MPO neural network to reduce the parameter count of LoRA.
-The input vector of LoRA is reshaped and fed into $\mathrm{MPO_{A}}$, where the last dimension of the output is output mpo $\mathrm{output}_{\mathrm{mpo_{A}}}$=128.
-This output is then transformed by a classical Multilayer Perceptron ($\mathrm{MLP_{A}}$) into output $\mathrm{output}_{\mathrm{MLP_{A}}}$=4, which corresponds to the number of qubits in the subsequent QNN.The output $\mathrm{output}_{\mathrm{MLP_{A}}}$ is input into the QNN, where it undergoes RY angle encoding and variational quantum state transformations.
-By exploiting the superposition and entanglement of quantum states, the QNN efficiently represents high-dimensional features in Hilbert space.
-Finally, the expectation values of PauliZ measurements on each qubit yield an output vector $\mathrm{output}_{\mathrm{QNN}}$ of the same size as $\mathrm{output}_{\mathrm{MLP_{A}}}$.
-The input to the next layer $\mathrm{MLP_{B}}$ is computed as: $\mathrm{input}_{\mathrm{MLP_{B}}}=\mathrm{output}_{\mathrm{QNN}}\odot\mathrm{output}_{\mathrm{MLP_{A}}}+\mathrm{output}_{\mathrm{MLP_{A}}}$, where $\mathrm{output}_{\mathrm{QNN}}\odot\mathrm{output}_{\mathrm{MLP_{A}}}$ corresponds to Equation (6).
-The $\mathrm{MLP_{B}}$ then linearly transforms input $\mathrm{output}_{\mathrm{MLP_{B}}}$ into $\mathrm{output}_{\mathrm{MLP_{B}}}$=128, which is fed into $\mathrm{MPO_{B}}$ to produce the final output of the LoRA layer based on QWTHN.
-All the above dimensions can be set as hyperparameters.
-We employ RY rotation angle encoding to reduce the number of encoding layers, thereby enhancing the training efficiency of the QNN and improving the scalability for real quantum hardware execution.
-By utilizing parameterized two-qubit gates CRZ (Controlled-Rotation-Z), we can adjust parameters to enhance the exploration of the state space.
+While MPO neural networks can generate non-harmonic functions, QNN excels at fitting truncated Fourier series [12, 15], by leveraging quantum state superposition and entanglement, QNN generate nonlinear features in high-dimensional Hilbert space.
+To address the limitations of classical neural networks in low-rank spaces, we linearly combine the classical output of the QNN, denoted as $O_{q}$, with the output of the classical neural network, denoted as $O_{c}$, $W_{q}$ and $W_{c}$ are the weights for the quantum neural network output and the classical neural network output, respectively, used to adjust the contributions of $O_{q}$ and $O_{c}$:
+$\tilde{O}=W_{q}O_{q}+W_{c}O_{c}.$ (10)
+The updated output $\tilde{O}$ retains the low-rank feature learning capability while incorporating the high-dimensional features extracted by the QNN, thereby enhancing the model’s ability to model complex nonlinear relationships.
+Through this linear operation, the QNN’s output $O_{q}$ provides additional nonlinear features to $\tilde{O}$.
+The elements in $\tilde{O}$ encompass both harmonic and non-harmonic features [16, 17, 18, 19], extracting harmonic features and complementing the classical neural network.
+As a result, the expressive power of the updated output $\tilde{O}$ is theoretically enhanced.
+Based on the aforementioned MPO and quantum hybrid network, we construct the final QTHA, as shown in Fig.
+1.
+QTHA uses MPO to reduce the number of parameters in LoRA.
+Specifically, the input vector of LoRA is reshaped and fed into $\mathrm{MPO_{A}}$.
+The output from $\mathrm{MPO_{A}}$ is then processed by a classical multilayer perceptron ($\mathrm{MLP_{A}}$), which transforms it into a representation corresponding to the number of qubits in the subsequent QNN.
+Within the QNN, the input is first encoded using RY angles, followed by a variational quantum circuit.
+These operations leverage quantum superposition and entanglement to efficiently encode high-dimensional features.
+After performing Pauli-Z measurements on each qubit, a vector of the same length as the output of $\mathrm{MLP_{A}}$ is generated.
+The input for the next layer, $\mathrm{MLP_{B}}$, is computed as a weighted combination of the QNN output and the output from $\mathrm{MLP_{A}}$.
+Subsequently, $\mathrm{MLP_{B}}$ applies a linear transformation to this combined input, producing a vector that is then passed to $\mathrm{MPO_{B}}$ to generate the final output.
+All dimensions within this architecture can be configured as hyperparameters for flexibility and optimization.
+By using parameterized two-qubit gates CRZ (Controlled-Rotation-Z), we can adjust parameters to enhance the exploration of the state space.
 Our design adopts a block structure, where each block consists of continuous nearest-neighbor interactions and one non-local interaction.
 This modular approach supports performance enhancement through stacking multiple layers, yet achieves near-saturation of expressive power with fewer layers, thus reducing optimization complexity.
-This design is particularly well-suited for current quantum computing hardware.
+Figure 1: Schematic diagram of the Quantum Tensor Hybrid Adapation (QTHA). QTHA dynamically adjusts feature weights through parameter tuning and outputs a combination of features from MPO and QNN.
 III Experiment
 III-A Datasets
-The datasets used in this study are publicly available, specifically the CPsyCoun[20] Chinese dataset, the R1-Distill-SFT[21] English dataset and the Chinese-DeepSeek-R1-Distill-data-110k[22].
-The CPsyCoun Chinese dataset is based on the open-source version available on the HuggingFace platform, released by the CAS-SIAT-XHai team.
-It is a psychological counseling dialogue dataset containing real-world Chinese dialogue records from counseling scenarios, covering nine common psychological issues such as depression, anxiety, and interpersonal relationships.
-The dataset spans seven classic theoretical schools in the field of psychological counseling, representing different theoretical frameworks in counseling practice, and is considered a representative branch of technical methods.
-The data has been anonymized and annotated with multi-turn dialogue roles between counselors and clients, comprising approximately 16,000 dialogue samples.
-Some data includes fine-grained labels such as counseling strategies and types of emotional support.
-This dataset is suitable for research on psychological counseling dialogue generation, sentiment analysis, and intervention strategy modeling, providing a benchmark for exploring the application of natural language processing technologies in the mental health domain.
-The R1-Distill-SFT dataset is one of the open-source datasets provided by the ModelScope community.
-It is a structured question bank aimed at K-12 mathematics education, covering areas such as arithmetic, algebra, geometry, and comparison, along with step-by-step solutions.
-Each data sample includes the question stem, blanks to be filled, multilingual labels (e.g., English and Chinese cn_k12), standardized answers (annotated using LaTeX mathematical symbols), and verification status.
-The dataset is suitable for model training (e.g., Supervised Fine-Tuning, SFT) and automated evaluation in educational scenarios.
-It has been filtered based on quality metrics, covering topics ranging from basic operations to medium complexity, and supports research on personalized learning and reasoning ability assessment.
-The R1-Distill-SFT dataset focuses on supporting tasks such as Knowledge Distillation and Supervised Fine-Tuning (SFT) for models.
-The Chinese-DeepSeek-R1-Distill-data-110k is one of the Chinese datasets provided by the ModelScope community.
-It is a Chinese instruction fine-tuning dataset generated by distilling DeepSeek-R1, containing approximately 110,000 high-quality multi-turn dialogue samples.
-This dataset is designed to support the fine-tuning and optimization of the DeepSeek-R1 series of models.
-It covers a wide range of domains, including general dialogue, knowledge-based question answering, code generation, and mathematical reasoning.
-The dataset is generated using distillation techniques from large language models and undergoes rigorous cleaning and deduplication processes.
-The data is structured into input-output pairs with contextual annotations, ensuring high diversity in instructions, strong logical coherence, and broad task coverage.
-It is particularly suitable for training dialogue models capable of understanding and executing complex instructions.
+The datasets used in this study are publicly available, specifically the CPsyCoun [20] Chinese dataset, the R1-Distill-SFT [21] English dataset and the Chinese-DeepSeek-R1-Distill-data-110k (CH-R1-Math) [22].
+CPsyCoun dataset is a Chinese psychological counseling dialogue dataset from HuggingFace (by CAS-SIAT-XHai), featuring anonymized, multi-turn dialogues (16K samples) across 9 common issues (e.g., depression, anxiety) and 7 counseling schools.
+Includes counselor-client role labels, some with fine-grained annotations (e.g., strategies, emotional support).
+Ideal for counseling dialogue generation, sentiment analysis, and intervention modeling in NLP for mental health.
+R1-Distill-SFT dataset is a K-12 math QA dataset with structured questions (arithmetic, algebra, geometry) and step-by-step solutions.
+Includes question stems, multilingual labels (e.g., English/Chinese), LaTeX-formatted answers, and verification status.
+Filtered for quality, it supports SFT, knowledge distillation, and automated evaluation in education.
+CH-R1-Math dataset is a Chinese instruction-tuning dataset (110K samples) from ModelScope, distilled from DeepSeek-R1.
+Covers multi-turn dialogues, QA, code generation, and math reasoning.
+Rigorously cleaned and deduplicated, it features diverse tasks, logical coherence, and contextual annotations.
+Optimized for fine-tuning models like DeepSeek-R1.
 III-B Evaluation Metrics
-In this experiment, Cross-Entropy Loss[23] is adopted as the evaluation metric for the model.
+In this experiment, cross-entropy loss[23] is adopted as the training and validation loss function for the model.
 Cross-entropy measures the difference between the predicted probability distribution of the model and the true label distribution, effectively reflecting the optimization level of classification tasks.
@@
 This loss function averages the prediction errors across all samples through the double summation operation, effectively penalizing misclassified predictions while ensuring comparability of losses across different batches of data.
-For text generation evaluation, we utilize BLEU-4[24] and ROUGE[25] scores. BLEU-4 calculates 4-gram precision with a brevity penalty BP to penalize short outputs:
-[…]
-$\text{BP}=\begin{cases}1&\text{if }c>r\\ e^{(1-r/c)}&\text{if }c\leq r\end{cases},$
-$\text{Precision}_{n}=\frac{\sum\text{Count}_{\text{clip}}(\text{n-gram})}{\sum\text{Count}_{\text{generated}}(\text{n-gram})},$
-  • BLEU-4: Bilingual Evaluation Understudy score for 4-gram precision,
-  • BP: Brevity Penalty to penalize short translations,
-  • $w_{n}$: Uniform weights with $w_{n}=\frac{1}{4}$,
-  • $\text{Precision}_{n}$: Modified n-gram precision for $n=1$ to $4$,
-  • $c$: Total length of candidate translations,
-  • $r$: Effective reference corpus length,
-[…]
-[…]
-The recall metric ROUGE-N quantifies the proportion of n-grams in the reference(s) text that are successfully reproduced in the generated text. Formally, it is computed as:
-[…]
-  • R: the set of reference text(s),
-[…]
-[…]
-The ROUGE-L metric evaluates text generation quality by measuring the alignment between a generated text and reference text(s) using the Longest Common Subsequence (LCS).
-Unlike n-gram-based metrics, ROUGE-L emphasizes semantic coherence and structural continuity by identifying the longest sequence of tokens that appear in both texts in the same order, though not necessarily contiguously.
-The metric is computed as a balanced F1-score derived from LCS-based recall ($R_{\text{LCS}}$) and precision ($P_{\text{LCS}}$):
-[…]
-$R_{\text{LCS}}=\frac{\text{Length of LCS(Ref,Gen)}}{\text{Total tokens in reference text(s)}},$
-$P_{\text{LCS}}=\frac{\text{Length of LCS(Ref,Gen)}}{\text{Total tokens in generated text}},$
-  • $R_{\text{LCS}}$: Recall of the Longest Common Subsequence,
-  • $P_{\text{LCS}}$: Precision of the Longest Common Subsequence,
-  • $\beta$: Controls the relative weight of recall ($R_{\text{LCS}}$) and precision ($P_{\text{LCS}}$). Typically, $\beta$=1 to prioritize harmonic mean (F1).
-III-C Based Model
-The baseline model for comparison is the LoRA algorithm with a rank of 4, fine-tuned based on the DeepSeek-R1-Distill-Qwen-7B (https://modelscope.cn/models/deepseek-ai/DeepSeek-R1-Distill-Qwen-7B) and Qwen2-7B-Instruct models.
-III-D Results
-This section primarily aims to verify whether the efficient fine-tuning algorithm described for the QWTHN can improve the performance of fine-tuned models while significantly reducing the number of parameters.
-This evaluation is based on the assumption that the QWTHN can enhance the expressive power of classical neural networks in low-rank spaces.
+Perplexity (PPL) is a common evaluation metric in natural language processing used to measure how well a language model predicts a text sequence.
+It reflects the model’s uncertainty when assigning probabilities to words in the test data.
+Lower perplexity indicates better performance.
+PPL is defined as:
+$\text{PPL}=\exp\left(-\frac{1}{N}\sum_{i=1}^{N}\log P(w_{i}|w_{<i})\right),$
+where $N$ denotes the total number of words/tokens in the test data and $P$ is the probability assigned by the model to the $i$-th word given prior context.
+A lower PPL means that the model assigns higher probabilities to the correct words, showing better alignment with the true data distribution.
+For the evaluation of text generation, we use BLEU-4 [24] and ROUGE [25].
+BLEU-4 assesses lexical precision through 4-gram matching between generated and reference texts, while ROUGE measures content coverage and semantic coherence using recall-oriented n-gram and sequence alignment metrics.
+This combination provides complementary perspectives on the quality of the generation.
+III-C Results
+Benchmark of various fine-tuning methods using DeepSeek-R1-Distill-Qwen-7B [26] and Qwen2-7B-Instruct [27] models as the base model.The ranks of LoRA [1] and QTHA are set to 4, and the decomposition factor of QuanTA [5] is set to 5.
+QTHA, LoRA and QuanTA are applied to the linear projection layers q_proj and v_proj.
+This section primarily aims to verify whether the efficient fine-tuning algorithm described for the QTHA can improve the performance of fine-tuned models while significantly reducing the number of parameters.
 The experiments were conducted using PyTorch [28] and PyVQNet [29, 30] with quantum circuit simulations.
-In this stage, noise effects in the quantum system are ignored, and the Cross-Entropy Loss is adopted as the optimization objective for classification tasks.
-As shown in Fig 3, the convergence curve of QWTHN exhibits a steeper descent rate during early training stages and achieves a lower stable loss value in training phase.
-Figure 3: Training loss comparison of QWTHN, QuanTA, and LoRA across varying sample sizes (N) and datasets.
-Subplots (a–c) illustrate results on the CPsyCounD dataset with N={300,600,1000}, while subplots (d–f) show corresponding analyses on the R1-Distill-SFT dataset.
-QWTHN consistently achieves lower training losses compared to baseline methods, demonstrating enhanced convergence efficiency under diverse experimental conditions.
-The datasets are extracted from CPsyCounD and R1-Distill-SFT with 300, 600, and 1000 samples, respectively.
+As shown in Fig.
+2, the convergence curve of QTHA exhibits a steeper descent rate during early training stages and achieves a lower stable loss value in training phase.
+Figure 2: Training loss comparison of QTHA, LoRA and QuanTA across varying sample sizes (N) and datasets.
+Subplots (a–b) illustrate results on the CPsyCounD dataset, while subplots (c–d) and (e–f) show corresponding analyses on the R1-Distill-SFT and CH-R1-Math dataset.
+QTHA consistently achieves lower training losses compared to baseline methods, demonstrating enhanced convergence efficiency under diverse experimental conditions.
+The datasets are extracted from CPsyCounD, R1-Distill-SFT, and CH-R1-Math, with 300, 600, 1000, 3000 respectively.
 The maximum sequence length max_seq_len for both training and prediction is set to 1024, with 10% of the data used as the validation set and the remaining 90% as the training set.
-For the Qwen2-7B-Instruct test set, the last 50 samples from CPsyCounD (with no overlap with the previous data) are used as the test set.
-Since DeepSeek-R1-Distill-Qwen-7B excels in code and mathematical capabilities, using metrics like BLEU-4 to compare the performance of QWTHN LoRA and MLP LoRA is not appropriate.
-Instead, we evaluate the test set performance using the following metrics:
+For the test set, 30 to 100 samples from the aforementioned three datasets are extracted as the test set.
+DeepSeek-R1-Distill-Qwen-7B exhibits excellent performance in code and mathematical reasoning tasks.
+we use the following metrics to perform a small-scale evaluation of the performance of the test set:
   • Strict Accuracy (SA): Defined as whether the output contains the correct answer in the answer segment.
@@
   • Answer Completeness (AC): Measures the completeness of the answer itself.
-As demonstrated in Table I, the proposed QWTHN achieves significantly higher training efficiency than classical LoRA under identical fine-tuning conditions for LLMs.
-When evaluated on chain-of-thought (CoT) reasoning tasks, QWTHN exhibits statistically significant improvements in domain-specific metrics—including Contextual Task Consistency (CTC), Answer Coherence (AC), Accuracy, and Strict Accuracy—while reducing trainable parameters by over 70% compared to LoRA.
-Further validation across benchmark datasets (CPsyCounD, R1-Distill-SFT, and Chinese-DeepSeek-R1-Distill-data-110k) confirms the robustness of QWTHN.
-As summarized in Table II, the framework demonstrates consistent gains in text generation quality, with absolute increases of 1–3 points in BLEU-4, ROUGE-1, ROUGE-2, and ROUGE-L metrics.
-These results underscore QWTHN’s ability to reconcile parameter efficiency with enhanced model performance, establishing it as a viable alternative to conventional low-rank adaptation paradigms.
-Table I: Cross-Architecture Performance Comparison by Dataset Scale
+As demonstrated in Table I, the proposed QTHA achieves significantly higher training efficiency than classical LoRA under identical fine-tuning conditions for LLMs.
+When evaluated on chain-of-thought (CoT) reasoning tasks, QTHA exhibits statistically significant improvements in domain-specific metrics—including Contextual Task Consistency (CTC), Answer Coherence (AC), Accuracy, and Strict Accuracy—while reducing trainable parameters by over 76% compared to LoRA.
+Further testing on benchmark datasets (CPsyCounD, R1-Distill-SFT, and CH-R1-Math) confirms the robustness of QTHA.
+As summarized in Table II and III, the framework demonstrates consistent improvements in text generation quality, with notable enhancements observed in the BLEU-4, ROUGE-1, ROUGE-2, ROUGE-L, PPL metrics.These results underscore QTHA’s ability to reconcile parameter efficiency with enhanced model performance, establishing it as a viable alternative to conventional low-rank adaptation paradigms.
+Table I: Cross-Architecture Performance Comparison by Dataset Scale.
 CPsyCounD (Training/Validation Loss)
-Model Small (300) Medium (600) Full (1000)
+Dataset Scale LoRA QuanTA QTHA
 Train Val Train Val Train Val
-MLP 1.805 1.606 1.657 1.514 1.577 1.447
-quanTA 1.647 1.512 1.565 1.478 1.520 1.437
-QWTHN 1.529 1.475 1.512 1.464 1.497 1.464
+300 1.805 1.606 1.647 1.512 1.484 1.434
+600 1.657 1.514 1.565 1.478 1.452 1.421
+1000 1.577 1.447 1.520 1.437 1.440 1.405
+3000 1.764 1.600 1.633 1.517 1.572 1.494
 R1-Distill-SFT (Training/Validation Loss)
-Model Small (300) Medium (600) Full (1000)
+Dataset Scale LoRA QuanTA QTHA
 Train Val Train Val Train Val
-MLP 0.179 0.182 0.172 0.158 0.169 0.160
-quanTA 0.167 0.173 0.164 0.159 0.163 0.159
-QWTHN 0.155 0.174 0.158 0.166 0.162 0.164
-Note. Parameter counts: MLP = 1.26M (100%), quanTA = 1.37M (108.7%), QWTHN = 0.30M (23.7%).
-Table II: Performance Metrics for CPsyCounD, R1-Distill-SFT and Chinese-DeepSeek-R1-Distill-data-110k
+300 0.179 0.182 0.173 0.178 0.160 0.169
+600 0.172 0.162 0.170 0.163 0.159 0.158
+1000 0.169 0.160 0.168 0.163 0.160 0.159
+3000 0.229 0.213 0.220 0.208 0.211 0.202
+CH-R1-Math (Training/Validation Loss)
+Dataset Scale LoRA QuanTA QTHA
+Train Val Train Val Train Val
+300 0.880 0.876 0.855 0.853 0.804 0.804
+600 0.829 0.788 0.821 0.791 0.776 0.767
+1000 0.800 0.765 0.804 0.781 0.773 0.763
+3000 0.852 0.819 0.830 0.803 0.789 0.762
+Note. Parameter counts: LoRA = 1.26M (100%), QuanTA = 0.73M (57.5%), QTHA = 0.30M (23.7%), additionally, the number of parameters trained based on R1-Distill-SFT is 0.57M(45.2%).
+Table II: Performance Metrics for CPsyCounD, R1-Distill-SFT and CH-R1-Math.
+Here, “Simulator” denotes the results of QTHA produced by a quantum circuit simulator; and “Wukong” produced by the superconducting quantum computer “Origin Wukong”.
 CPsyCounD
 Model BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
-MLP 13.459 34.540 13.109 30.627
-quanTA 8.650 29.463 9.780 25.539
-QWTHN 14.593 37.345 14.023 32.32
-R1-Distill-SFT
-Model BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
-MLP 59.635 64.195 42.627 43.326
-quanTA 59.096 63.512 41.992 42.227
-QWTHN 61.615 66.087 45.294 45.915
-Chinese-DeepSeek-R1-Distill-data-110k
+LoRA 14.823 39.560 17.216 34.239
+QuanTA 13.368 37.860 15.257 33.705
+Simulator 16.043 40.638 17.614 35.892
+Wukong 18.056 42.668 20.322 37.172
+CH-R1-Math
 Model CTC AC Accuracy Strict Accuracy
-MLP 78% 68% 82% 68%
-quanTA 84% 74% 82% 68%
-QWTHN 96% 94% 86% 82%
-III-E Quantum Running
+LoRA 90 % 83 % 97 % 83 %
+QuanTA 97 % 93 % 97 % 93 %
+Simulator 100 % 100 % 93 % 93 %
+Wukong 100 % 100 % 100 % 100 %
+Table III: PPL for CPsyCounD, R1-Distill-SFT and CH-R1-Math. “Simulator” or “Wukong” means QTHA executed on a quantum circuit simulator or “Origin Wukong”, respectively.
+Model PPL(CPsyCounD) PPL(CH-R1-Math) PPL(R1-Distill-SFT)
+LoRA 5.5785 2.5235 1.3355
+QuanTA 5.0758 2.4675 1.3306
+Simulator 4.8350 2.3556 1.3241
+Wukong 4.8325 2.3550 1.3237
+III-D Implementation on quantum hardware
+QTHA demonstrates excellent performance on noisy quantum computers, while the output of LLMs is inherently a probability distribution matrix, its impact on the final result is always constrained within a specific threshold range, naturally exhibiting stochastic characteristics.
+The inherent random noise in quantum systems (such as decoherence and gate operation errors) may enhance reasoning robustness through the following mechanisms: in the probability matrix output by LLMs, only when the probability of a key token exceeds a preset threshold will it significantly influence decision-making.
+This implies that minor perturbations introduced by noise may be filtered out by these thresholds, thereby maintaining result stability.
+The probabilistic nature of quantum systems (e.g., superposition state collapse) and the probabilistic output of LLMs share mathematical similarities, and noise may implicitly calibrate the probability distribution to improve reasoning consistency.
 We have constructed a variational quantum circuit module incorporating the Origin Quantum Cloud (QCloudService) in our model architecture.
-This module utilizes the quantum chip backend codenamed "origin_wukong" to transform user-defined variational quantum circuits (including input data and trainable parameters) into quantum circuit intermediate representations (originIR) compatible with the QPanda framework[29].
-Given that Origin Quantum chips exclusively support Z-axis observable measurements, basis transformation operations are implemented to convert user-defined non-Z observables into equivalent Z-axis measurements.
+This module uses a quantum cloud service with a superconducting quantum computer backend named “Origin Wukong”, transforming user-defined variational quantum circuits (including input data and trainable parameters) into quantum circuit intermediate representations (OriginIR) [31].
+Given that the backend exclusively support Z-axis observable measurements, basis transformation operations are implemented to convert user-defined non-Z observables into equivalent Z-axis measurements.
 Specifically, each Pauli operator measurement generates a distinct quantum circuit.
@@
 Consequently, the total number of quantum circuits executed per batch on the quantum hardware is governed by the relationship:
-$\mathrm{T=B\times Q},$
+$\mathrm{T=B\times Q},$ (11)
 where B denotes the batch size and Q represents the number of qubits in the system.
@@
 Finally, measurement results are reformatted into machine learning-compatible tensors through concatenation and reshaping operations.
-These processed outputs can be directly utilized by subsequent neural network modules, establishing seamless integration between quantum computations and classical machine learning pipelines.
-III-F Discussion
-The parameter count of the Quantum Tensor Hybrid Network (QWTHN), which combines quantum tensor networks and quantum neural networks, is lower than that of classical LoRA, while the training loss is significantly reduced.
-This leads to improved parameter efficiency.
-The Matrix Product Operator (MPO) decomposition introduces the concept of rank, effectively decomposing high-dimensional linear transformations into a series of low-dimensional matrix products.
-This decomposition significantly reduces the number of model parameters, especially when handling high-dimensional inputs and outputs.
-Traditional fully connected layers flatten the input into vectors, losing the dimensional structure of the original data.
-In contrast, MPO neural networks preserve this structural information to some extent because they operate on the dimensions of the original tensor.
-This structured approach may enable MPO to learn and represent multidimensional relationships and patterns in the data more effectively, such as correlations between dimensions and hierarchical structures.
-By leveraging the classical output of the QNN to perform element-wise multiplication with the output of the classical network, the expressive power of classical neural networks in low-rank spaces is enhanced.
-Quantum neural networks process input data using quantum properties (e.g., superposition, entanglement, and parallel computation), potentially generating high-dimensional or nonlinear features.
-Even though the final output of the QNN is classical data (e.g., probability distributions after measurement), the underlying quantum state evolution process can still extract information that is difficult to obtain using classical methods.
-Currently, the scale of dataset experiments is limited by the size of quantum circuits.
-Although it can be seen that training loss on a larger dataset of 1000 samples still performs better.
-Future work will need to focus on developing fine-tuning strategies for larger datasets and even exploring large-scale quantum pre-training schemes.
-The current integration of MPO (Matrix Product Operator) and VQC (Variational Quantum Circuit) involves hard connections between input and output interactions.
-However, theoretically, all weight matrices could be replaced with variational quantum circuits, which would enable the model to capture higher degrees of correlations far beyond what classical weight matrices can represent.
-Weight matrix W can be decomposed into an MPO with a given bond dimension, and this MPO can be further disentangled using quantum circuits, with the remaining operators also expressed in MPO format [30].
-This is one feasible technical direction, and we will explore more potential implementations of large-scale quantum models.
+These processed outputs can be directly processed by subsequent neural network modules, establishing seamless integration between quantum computations and classical machine learning pipelines.
 IV Conclusion
-This paper implements a quantum-weighted tensor hybrid fine-tuning algorithm for large models, leveraging quantum computing principles to significantly reduce parameter redundancy and enhance parameter efficiency.
-It validates the potential of quantum computing to alleviate the computational bottleneck of large models, achieving a novel integration of quantum computing and large-scale models.
-At the current stage of quantum computer development, this work not only demonstrates the efficient adaptation of quantum resources to billion-parameter models for lightweight optimization but also validates a practical pathway for quantum hardware driven by large model tasks.
-This lays the first engineerable technical foundation for future quantum-enhanced Artificial General Intelligence (AGI) systems.
+This paper introduces QTHA, a quantum tensor hybrid adaptation framework for fine-tuning LLMs, achieving significant breakthroughs in parameter reduction and performance enhancement by integrating quantum tensor networks with quantum neural networks.
+Leveraging MPO decomposition, the framework transforms high-dimensional linear transformations into low-dimensional matrix product sequences via rank constraints, drastically reducing parameter counts while preserving tensor structures to effectively capture multi-dimensional data correlations.
+Additionally, the QNN generates high-dimensional nonlinear features unattainable by classical methods, and its fusion with the MPO network significantly enhances representation power in low-rank spaces.
+Experimental results demonstrate superior training loss reduction, even with limited datasets of 3,000 samples, highlighting quantum computing’s potential to address computational bottlenecks in LLMs.
+Looking forward, the research will explore large-scale quantum pre-training schemes and fully quantum weight matrix reconstruction using variational quantum circuits to capture higher-order correlations.
+This work not only achieves lightweight optimization for billion-parameter models but also establishes a practical paradigm where LLM-driven tasks advance quantum hardware development.
+It lays the foundation for future quantum-enhanced Artificial General Intelligence systems, marking a key step toward scalable and engineerable quantum AI solutions.